OpenStudy (anonymous):

–2(x – 1) = 16 –2(x – 1) = 16 @Mathematics

OpenStudy (diyadiya):

-2x+2=16 -2x=14 x=-7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

x = -7

OpenStudy (anonymous):

-7